# V-FLUTE: Visual Figurative Language Understanding with Textual Explanations

### v1

@@ D.1.1 One-shot Prompt for generating explanations
D.2 IRFL
D.2.1 Zero-shot Prompt for generating explanations
@@
D.3 MuSE
D.3.1 Few-shot Prompt for generating opposite claims
@@
D.3.2 Zero-shot Prompt for Rephrasing
@@
D.4 MemeCap
D.4.1 Few-shot Prompt for generating entailing claims
We describe our few-shot prompts given to an LLM (gpt-4-0613) for generating entailing claims as part of the pipeline. Refer to Table 13 for the detailed prompt.
D.4.2 Zero-shot Prompt for validating the entailing claims
@@
D.4.3 Few-shot Prompt for generating opposite claims
@@
D.4.4 Zero-shot Prompt for generating explanations
@@
Appendix E Model Taxonomy
@@
Appendix F By-Phenomenon Performance
@@
Appendix G Annotator Recruitment and Compensation
All recruited annotators have significant background in figurative language understanding (having formal educational background in linguistics or literature). Annotators were recruited through Upwork platform. All are on fluent or native/bilingual level in English. All are fairly compensated with USD $20 to $25 per hour with self-reported time needed to complete the tasks.

### v2

@@ F.1.1 One-shot Prompt for generating explanations
F.2 IRFL
F.2.1 Zero-shot Prompt for generating explanations
@@
F.3 MuSE
F.3.1 Few-shot Prompt for generating opposite claims
@@
F.3.2 Zero-shot Prompt for Rephrasing
@@
F.4 MemeCap
F.4.1 Few-shot Prompt for generating entailing claims
We describe our few-shot prompts given to an LLM (gpt-4-0613) for generating entailing captions as part of the pipeline. Refer to Table 12 for the detailed prompt.
F.4.2 Zero-shot Prompt for validating the entailing captions
@@
F.4.3 Few-shot Prompt for generating opposite claims
@@
F.4.4 Zero-shot Prompt for generating explanations
@@
Appendix G Model Taxonomy
@@
Appendix H Multimodal Structured Chain-of-Thought Performance
In addition to zero-shot testing, we also test these models using Compositional Chain-of-Thought Prompting proposed by Mitra et al. (2023). The method prompts the model zero-shot to generate a scene graph in JSON format and then utilizes that scene graph in another prompt to answer the relevant question. We refer to these models as LLaVA-ZS-7B-SG and LLaVA-ZS-34B-SG for the 7B and 34B LLaVA configurations described above.
Scene graph prompting and few-shot prompting improves performance on the figurative visual entailment task.
Observing the results in Table 3, we can see that the multimodal few-shot prompting and scene graph prompting, having demonstrated their effectiveness for literal inputs, also show improved performance on the figurative visual entailment task. However, the explanations generated by SG-models tend to overly focus on the contents of the scene graph rather than the underlying figurative phenomena, possibly causing a decrease in explanation score.
Appendix I Additional Models
In addition to the LLaVA architecture, we conduct experiments with the Instruct-BLIP model Dai et al. (2024), specifically, the Instruct-BLIP-Vicuna-7B version. As can be seen in Table 16, InstructBLIP shows a weaker performance compared to LLaVA-7B, especially in explanation quality (4.14 F1@53 for InstructBLIP while 35.56 for LLaVA-7B-ZS, and 2.07 F1@60 while 18.38 for LLaVA as can be seen in Table 3). It struggled to generate scene graph descriptions, unlike LLaVA-7B. Despite extensive instruction-tuning, it performed below a random baseline in our figurative entailment task (F1@0: 43.37).
| Model Name | f1@0 | f1@53 | f1@60 |
| InstructBlip-7B-ZS | 43.37 | 4.14 | 2.07 |
| InstructBlip-7B-SG | 38.03 | 4.15 | 1.38 |
We also experimented with a state-of-the-art multimodal model GPT-4o that was released after our dataset was created. As expected, the results are better than those of GPT-4 due to improvements in the multimodal processing of GPT-4o. However, the F1@53 and F1@60 scores suggest there could still be improvements in explanation quality. Compared to the 7B fine-tuned LLaVA model, the zero-shot GPT-4o still underperforms the fine-tuned models in terms of F1@53 and is comparable in terms of F1@0. GPT-4o in the few-shot scenario (5 example) shows better results than the fine-tune model. These results can add to the discussion in our field between smaller open-source models and bigger and proprietary models in terms of performance accuracy and capabilities.
| Model Name | f1@0 | f1@53 | f1@60 |
| GPT-4o | 75.41 | 60.97 | 37.20 |
| 5-shot | 79.42 | 69.35 | 56.31 |
Appendix J By-Phenomenon Performance
@@ Appendix I Additional Models
Appendix K Human Baseline
| Phenomenon | Dataset |
| Metaphor /Similes | HAIVMET | 78.84 | 81.25 | ||||
94.36 | 77.78 | |||||
| Idioms |
89.26 | 49.74 | ||||
| Sarcasm | MuSE | 68.89 | 85.42 | ||||
| Humor | MemeCap | 100.0 | 78.03 | ||||
| NYCartoons | 71.43 | 47.83 | |||||
| Overall | 89.09 | 77.26 | |||||
Appendix L How Well Does the Explanation Score Predict Human Judgment on Adequacy?
We explore whether the proposed explanation score can capture human judgement of explanation adequacy. We collect all instances from Section 5 where both annotators agreed on the adequacy judgement for the explanation. We evaluate if the explanation score described in Section 4.2 can act as a good predictor for the human adequacy judgment. We find that the area under the Precision-Recall curve is 0.79, and the maximum F1 score is 0.77, obtainable at the explanation score threshold of 0.53. Hence, we use this threshold to report the results in Table 3. We also use the threshold of 0.6 since it maximizes F1 such that both precision and recall are above 0.75.
Appendix M How Do Models Perform When Only Predicting the Label?
In our experiments, we found that predicting only the label improves accuracy compared to predicting label and explanation (this is expected and observed in other work on textual explanations such as e-SNLI Camburu et al. (2018)). However, these predictions are less reliable since they could be due to spurious correlations (which is why we require the model to generate textual explanations). We also found when fine-tuning the model in a multi-task fashion with explanations (i.e., two tasks, one of generating explanations and one of predicting the label), the accuracy improves compared to when fine-tuning only for the prediction task (F1 score of 80.85 vs. 83.26, ), in line with previous findings by Hsieh et al. (2023).
Appendix N Annotation Interfaces
We provide the annotation interfaces below for HAIVMET (Figure 12), IRFL (Figure 13), MemeCap (Figure 14) and MuSE (Figure 15). In addition, instructions were explained in more detail to the annotators via chat on Upwork, and any of their doubts and questions were answered.
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/vismet_Ex2.png)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/irfl_example_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/muse_ex_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/memecap_ex_resized.jpeg)
![[Uncaptioned image]](extracted/5929537/figs/dataset_exs/ny_ex_resized.jpeg)